### v2

# Configurable DOA Estimation using Incremental Learning

Abstract
This study introduces a progressive neural network (PNN) model for direction of arrival (DOA) estimation, DOA-PNN, addressing the challenge due to catastrophic forgetting in adapting dynamic acoustic environments. While traditional methods such as GCC, MUSIC, and SRP-PHAT are effective in static settings, they perform worse in noisy, reverberant conditions. Deep learning models, particularly CNNs, offer improvements but struggle with a mismatch configuration between the training and inference phases. The proposed DOA-PNN overcomes these limitations by incorporating task incremental learning of continual learning, allowing for adaptation across varying acoustic scenarios with less forgetting of previously learned knowledge. Featuring task-specific sub-networks and a scaling mechanism, DOA-PNN efficiently manages parameter growth, ensuring high performance across incremental microphone configurations. We study DOA-PNN on a simulated data under various mic distance based microphone settings. The studies reveal its capability to maintain performance with minimal parameter increase, presenting an efficient solution for DOA estimation.
Index Terms:
DOA Estimation, Sound Source Localization, Incremental Learning, Configurable Model
I Introduction
Sound source localization (SSL) [1] aims to identify the origins of sound sources by analyzing signals captured through an array of microphones. At the cornerstone of SSL lies the direction of arrival (DOA) estimation, a process crucial for calculating the angles at which sounds reach the microphone array. This spatial information is necessary for speech-related applications [2, 3, 4, 5]. These applications rely on the precise localization of sound sources to significantly improve their performance and functionality. Consequently, the development of more sophisticated DOA estimation techniques has become a key focus within the research community to achieve greater robustness and adaptability in real-world settings.
Classical signal processing techniques like generalized cross-correlation (GCC) [6], multiple signal classification (MUSIC) [7], and steered response power (SRP) based methods, particularly SRP-PHAT [8], which is obtained from the SRP by applying phase transform (PHAT) whitening. They are available to adapt to different microphone array configurations. However, their effectiveness decreases in noisy and reverberant environments compared with deep learning solutions.
A significant milestone in this exploration is the adoption of deep learning architectures such as convolutional neural networks (CNN) [9, 10, 11, 12, 13]. However, deep learning models used for DOA estimation face a unique challenge: the misalignment between training and testing environments, especially concerning the microphone array’s configuration. Such mismatches precipitate a decline in model efficacy, necessitating re-training for new configurations. Re-training always requires extra time and computation resources. This highlights an essential area for ongoing deep learning-based DOA estimation research, to enhance model accuracy across diverse acoustic settings. Although some methods [14, 15] adapt to the variable array setups relying on specific spatial features, they do not address more complex acoustic configurations like the difference absorption and reverberation time. Along the diverse configurations, one fixed method faces forgetting because the new knowledge overrides the existing knowledge.
Various research recently incorporated continual learning (CL) [16] to continuously learn new knowledge while retaining previously learned knowledge in audio classification [17, 18]. Task Incremental Learning (TIL)  [19] is a subset of CL methods. The TIL approach enables the model to perform new tasks efficiently with the same number of classes in each task without losing prior performance. Inspired by [20, 21, 22], we proposed to use TIL to gradually learn new information for DOA estimation using simple, widely used features like the short-time Fourier transform (STFT), all while keeping the valuable knowledge it has already gained from the existing acoustic environments and configurations.
In this research, we introduce DOA-PNN, a novel progressive neural network based on the TIL method for DOA estimation, crafted to solve the issue of adaptation from source to target configurations. The core of DOA-PNN’s architecture is task-specific sub-networks , each engineered to memorize knowledge from previous configurations, enabling smooth transitions across diverse acoustic environments while maintaining accuracy on previous tasks. A scaling mechanism is integrated into these sub-networks to limit parameter growth by reducing the extra layers. Contrary to traditional finetuning methods, DOA-PNN utilizes a collection of pre-trained sub-networks, establishing lateral connections that extract the crucial features for new tasks, thereby enriching feature presentation. To the best of our knowledge, this is the first work to apply continual learning techniques to adapt DOA estimation to various acoustic configurations. Demonstrated through extensive testing on a simulated dataset for DOA estimation, DOA-PNN exhibits competitive performance on unfamiliar configurations with less parameter increase and no reliance on previous data buffers.
Figure 1: (a) Architecture of SSNet (b) Proposed DOA-PNN framework.
II Proposed Method
II-A Continual learning for DOA estimation
In this work, applying continual learning to DOA estimation involves receiving a series of $T$ distinct tasks (under different acoustic configurations) in sequence and optimizing the performance across all tasks with less negative impact of catastrophic forgetting. For each task $\tau_{t}$, with $t\leq T$, we handle a set of training data consisting of pairs $(x_{t},y_{t})$, where $x_{t}$ represents the acoustic signals and $y_{t}$ are the corresponding DOA labels, sampled from a distribution $D^{T}$.
Our goal is to minimize the cumulative loss $L_{tot}$, which aggregates the sub-losses $L_{doa}$ for all $T$ tasks, formulated as:
$L_{tot}=\sum_{t=0}^{T}E_{(x_{t},y_{t})\sim D^{T}}[L_{doa}(F_{t}(x_{t};\mu_{t}),y_{t})]$ (1)
where $L_{doa}$ represents a classification loss function, such as mean squared error (MSE) loss, and $F_{t}(x_{t};\mu_{t})$ is the DOA estimation model parameterized by $\mu_{t}$. During the training, we aim to find the optimal parameters $\mu_{t}^{\prime}$ that perform well across all $T$ tasks. This is challenging since parameters learned from previous task $\tau_{t}$ are prone to be overwritten when learning the subsequent task $\tau_{t+1}$, leading to catastrophic forgetting.
II-B Network backbone for DOA estimation
We adopt a spatial spectrum prediction module using the spatial spectrum network (SSNet) [23, 24, 25] for DOA estimation. We consider 2 microphone linear arrays as an instance in this section, although our proposed method can be employed for any number of microphone positions. This system learns complex relationships from audio data to predict, where the sound could be coming from at every degree in a half-circle of the linear microphone array. To perform this, we combine sound segments from different microphones and transform them into complex STFT. We consider the real as well as the imaginary part of the complex STFT instead of the phase and power and then combine this data to form a feature representation $v$. This $v$ feature is then used as the input to SSNet for predicting the spatial spectrum $\hat{p}$. The network structure of SSNet is shown in Fig. 1 (a) that depicts SSNet includes layers to simplify and analyze sound features. It uses two convolutional layers to reduce frequency details and five residual blocks to pick up intricate sound features. Then, it has a layer that focuses on the DOA estimation and rearranges the sound information in preparation for the final step. The result spatial spectrum is a 181-point map.
Following the work in [25], we use a probability-based method to determine the likelihood of a sound coming from any of the 180 directions. This part of SSNet assigns a probability to each direction based on how closely it matches the ground truth DOA. Specifically, each element of the encoded $181$-dimensional vector $p(\theta_{i})$ is assigned to a particular azimuth direction $\theta_{i}\in\{0^{\circ},1^{\circ},\dots,180^{\circ}\}$. The values of the vector follow a Gaussian distribution that maximizes at the ground truth DOA, which is defined as follows:
$p(\theta_{i})=\begin{cases}\exp\left(-\frac{d(\theta_{i},\theta_{y})^{2}}{2\sigma^{2}}\right),&\text{if }|\Theta|=0\\ 1,&\text{if }|\Theta|=1\end{cases}$ (2)
where $|\Theta|$ is the number of sources, $\theta^{\prime}$ is the ground truth DOA of one source, $\sigma$ is a predefined constant that controls the width of the Gaussian function and $d(\dots,\dots)$ denotes the angular distance. We then train SSNet by computing the MSE loss between the segment-wise output spatial spectrum from the model and the ground truth.
II-C Proposed DOA-PNN
Drawing on the concept of progressive neural networks, we introduce a novel progressive continual learning approach specifically designed for deep DOA estimation, depicted in Fig. 1(b). Our method, termed DOA-PNN, is composed of two principal elements: a task-specific sub-network creator for individual tasks and a lateral connection set bridging a suite of sub-networks. To manage the model size, we’ve introduced a gap-aware layer scaling mechanism in DOA-PNN, ensuring the parameter count remains controlled.
A task-specific sub-network creator is designed to generate the $t^{{\text{th}}}$ sub-network for each new task $\tau_{t}$. When a new acoustic configuration DOA estimation $\tau_{t}$ comes, the creator adds a sub-network with random initialization which is only the feature extractor part of SSNet. All sub-networks share the common memory which is the classification head of SSNet for a fixed class number 181.
A lateral connection set serve a pivotal role. When the $t^{{\text{th}}}$ sub-network processes input data, it not only forwards its output to the classification layer but also receives output features from all previously established sub-networks. To integrate these multiple outputs effectively, each sub-network is equipped with an adapter—a 1$\times$1 convolutional layer—tailored to its current context. This 1$\times$1 convolutional adapter transforms the combined features into the final output, thereby enriching the feature used for classification with the accumulated knowledge from the predecessors.
A gap-aware layer scaling mechanism adjusts the complexity of the SSNet feature extractor’s residual blocks based on a consistent parameter reflecting the minimum angle error tolerance. If the smallest acceptable error is 5 degrees, we streamline the residual blocks from five down to one. This manual scaling ensures the network remains efficient and tailored to the precision requirements of the DOA estimation.
In this architecture, for each new learning task labeled as $\tau_{t}$, we compute the real and imaginary parts of the complex STFT of audio segments to serve as input data. Concurrently, the DOA-PNN model preserves the integrity of the previously trained $t-1$ sub-networks by freezing their parameters, preventing interference with newly acquired knowledge. The creator then creates a fresh $t^{{\text{th}}}$ sub-network dedicated to the current learning task, leveraging both the STFT features and the task metadata—such as task id—to discern previously unidentified acoustic configurations. What sets the $t^{{\text{th}}}$ sub-network apart is its ability to access the shared memory of its predecessors through lateral connections, integrating all previous knowledge. When it comes to inference application, the DOA-PNN model selects the appropriate $t^{{\text{th}}}$ sub-network based on the specific task at hand for accurate and reliable evaluation. This approach ensures that our model remains adaptable and quick to respond in diverse acoustic settings, showcasing the potential of progressive continual learning strategy in the real-world environments of DOA estimation.
III Experiments
TABLE I: Parameter settings of simulated audio
Room Size [L, W, H] Absorption Tmax Num_images
[10,8,5], [5,4,3], [4,2,2] 0.2-0.8 0.2-0.6 [10,10,10]
III-A Dataset
We use the LibriSpeech [26] train-clean-100 subset, which comprises high-fidelity, single-channel speech recordings sampled at 16 kHz. It includes a diverse set of English speakers, selected for uniform audio segment lengths, ensuring homogeneity in data processing. From this dataset, we have curated a selection of 100 speakers out of the original 250. This selection balances gender representation with 47 male and 53 female speakers and encompasses a total of 1,511 audio files, amounting to approximately 4.96 hours of speech data.
We enhance the original single-channel speech data by creating two-channel versions with gpuRIR [27], arranging two linear microphones at distances ranging from 0.05 to 0.09 meters. The gpuRIR is a free and open-source Python library for room impulse response (RIR) simulation using the image source method with GPU acceleration. The details of the simulated room environments are specified in Table I. To ensure sound sources are well-distributed across the room, we first simulate the microphone center anywhere inside the room. Then we randomly choose a distance such that the source is inside the room to finally simulate the RIR. Based on the RIR, we generate the mixture data for each original single-channel speech data every 5 degrees from 0 to 180 degrees, which expands the audio files 37 times.
III-B Metrics
For evaluating our models, we utilize two primary metrics: mean absolute error (MAE) and accuracy (ACC). MAE measures the average magnitude of the errors in a set of predictions, without considering their direction. On the other hand, ACC is used to determine the percentage of predictions that are exactly correct or within a certain degree of tolerance. We assess accuracy at three levels of tolerance: Within a 5-degree, 10-degree, and 15-degree error margin. These tiers of tolerance allow us to understand not only the model’s precision but also its practical effectiveness.
(a) ACC ±$\pm 5$
(b) ACC ±$\pm 10$
(c) ACC ±$\pm 15$
Figure 2: Comparative inference performance of various DOA estimation methods across different microphone spacings, showing ACC within three tolerances.
III-C Incremental learning setting
In our incremental learning approach, we divide the dataset into five parts according to the varying distances between the microphones, with each partition representing a unique task. This allows us to treat the problem as a series of incremental learning tasks, each corresponding to a microphone spacing. We compare our approach with several baseline methods:
  • Finetune: Starting with a model pre-trained on the initial task, we sequentially fine-tune it on each subsequent task. This approach adapts the model incrementally, but may be susceptible to forget previously learned information when moving to new tasks each time.
  • Joint: A single model is trained on the entire dataset, encompassing all microphone spacings together. This method aims to learn a general representation that performs well across all the tasks.
  • Ensemble: This involves training five separate models, each on a different subset of the data, with each model specializing in a particular microphone spacing.
III-D Implementation details
For training each task within our DOA-PNN framework, we consistently apply 100 epochs with an early stopping criterion set to activate after 20 epochs if there’s no improvement in performance. We employ the AdamW [28] optimizer, a variant of the Adam optimizer that includes weight decay for regularization, with a learning rate of 0.001. Additionally, we apply a learning rate scheduler, specifically StepLR, to adjust the learning rate during training.
For the STFT computations, we set the frame size to 32 milliseconds and the hop size to 10 milliseconds. The frequency range for the STFT is confined between a minimum of 100 Hz and a maximum of 8000 Hz, using an n_fft value of 512, which determines the length of the windowed signal after padding with zeros. This STFT configuration is designed to capture a wide range of audio frequencies suitable for accurately determining the DOA. Gaussian function constant $\sigma$ we set as 8 in our experiments.
TABLE II: Performance comparison across different DOA estimation methods, highlighting MAE, ACC within various degree tolerances, and model parameters.
Methods MAE ACC(±$\pm 5$) ACC(±$\pm 10$) ACC(±$\pm 15$) Params
Finetune 17.8 33.9 48.0 58.4 0.9M
Joint 11.9 54.0 66.7 74.3 0.9M
Ensemble 7.6 78.6 84.8 87.5 4.8M
Proposed DOA-PNN 10.1 72.9 82.7 86.1 1.6M
IV Results and Discussions
We test the methods for DOA estimation across varying microphone spacings as Fig. 2, we can observe how each method adapts to the acoustic challenge. The finetune method incrementally trained from 5cm to 9cm only performs well for the recent task 9 cm indicating a forgetting across previously learned configurations. On looking to the performance for joint training, we find that it is more resilience compared to finetune, yet it does not maintain high performance as spacing widens. It is possible that the single-model approach is not able to capture the nuances needed for localization at variable spacings. The ensemble method demonstrates a robust performance across different spacings as the upper bound. It benefits from training distinct models for each spacing scenario.
Our proposed DOA-PNN shows a promising ability to sustain higher accuracy levels, particularly within the most challenging ±$\pm 5$-degree margin. It implies that our progressive learning strategy, which builds upon knowledge from previous configurations transfers well to new spacing scenarios. The proposed DOA-PNN offers a balance between accuracy and model complexity, suggesting it to be an efficient choice for real-world applications where microphone spacing can vary.
On examining the comparative results of different methods for DOA estimation as given in Table II, we find each method has its strengths and limitations. The finetune approach, while the most straightforward in terms of adapting pre-existing models to new tasks, seems to lag in precision. It presents the highest MAE, suggesting that it might not preserve the nuances of DOA estimation in different microphone spacing as effectively as the other methods. The joint training method improves upon finetune significantly, indicating that learning from a collective dataset of all tasks offers a better understanding of the DOA estimation problem. However, this method still does not reach the top performance in accuracy, especially in ±$\pm 5$-degree range. Ensemble, which employs a separate model for each condition, stands out in accuracy. It’s clear that specializing in individual conditions allows for a finer understanding of each specific DOA estimation task, but this comes at the cost of a larger overall model size, as reflected by the parameter count.
Our Proposed DOA-PNN method maintains a balance between precision and efficiency. It achieves comparable accuracy, nearly matching the best-performing method in the ±$\pm 15$-degree range and outperforming it within the ±$\pm 10$-degree margin. It achieves this while maintaining a moderate model size. This indicates that our progressive approach, which incorporates learning from successive tasks, effectively consolidates knowledge without increasing complexity.
We foresee several avenues for future work. While we currently focus on two-microphone linear arrays to estimate the azimuth angle, we plan to expand our research to encompass more complex acoustic configurations and to address elevation angle estimation for a full spatial representation [29]. Moreover, integrating noise robustness factors into DOA estimation is a critical next step, acknowledging that real-world applications often operate in acoustically challenging environments [30]. These enhancements will aim to bolster the practicality of DOA-PNN, providing comprehensive solutions for 3D sound localization in noisy, real-life settings.
V Conclusions
In this study, we presented DOA-PNN, an approach that advances the direction of arrival estimation within continual learning. DOA-PNN synergizes task-specific network instantiation with progressive learning principles, addressing the diversity in microphone configurations with a novel scaling mechanism and lateral connections for knowledge retention. Demonstrated across simulated acoustic settings, DOA-PNN consistently achieved high accuracy, underscoring its capacity for real-world applicability in dynamic auditory scenes. The model’s efficiency and adaptability suggest its potential as a foundational framework for future research in sound localization and acoustic signal processing, encouraging the development of sophisticated, scalable audio analysis technologies.